THE NEAR-INFRARED SPECTRUM OF CH_{2}^{+}

\underline{CHRISTOPHER F. NEESE}, JENNIFER L. GOTTFRIED, and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, The University of Chicago, Chicago IL, 60637, USA.

The molecular ion CH_{2}^{+} is of special theoretical interest because it is both quasi-linear and exhibits a strong Renner-Teller interaction between its ground and first-excited electronic states. At linearity, the ground state is a ^{2}Π_{u} state that splits into X̃^{2}A_{1} and Ã^{2}B_{2} states as the molecule bends. The Ã state is linear, while the X̃ is quasi-linear with a barrier to linearity of only 1089 cm^{−1}. Thus, only the ground vibrational state is bound by the barrier to linearity.

The spectrum of CH_{2}^{+} in the region 11,000–13,000 cm^{−1} has been recorded with our Ti:sapphire laser spectrometer. This spectrometer couples velocity modulation with heterodyne detection for near shot-noise-limited sensitivity. Since our initial letter on this spectrum,^{a} we have selectively rescanned portions of this spectrum with improved signal-to-noise. As a result, we have been able to assign the Ã(0, 3, 0)^{3} ← X̃(0, 0, 0)^{2} band and detect the X̃(0, 9, 0)^{2} ← X̃(0, 0, 0)^{1} band, whose assignment is in progress. A more detailed analysis of the entire spectrum, including spin splitting, is underway.

aJ. L. Gottfried and T. Oka, J. Chem. Phys. 121, 11527 (2004).